### v2

# On a question of Erdős and Gimbel on the cochromatic number

Abstract
In this note, we show that the difference between the chromatic and the cochromatic number of the random graph $G_{n,1/2}$ is not whp bounded by $n^{1/2-o(1)}$, addressing a question of Erdős and Gimbel.
1 Introduction
The cochromatic number $\zeta(G)$ of a graph $G$ is the minimum number of colours needed for a vertex colouring where every colour class is either an independent set or a clique. If $\chi(G)$ denotes the usual chromatic number, then clearly $\zeta(G)\leqslant\chi(G)$. Using classical methods, it is not hard to show that for the random graph $G_{n,1/2}$, with high probability [^1] As usual, we say that a sequence $(E_{n})_{n\geqslant 0}$ of events holds with high probability (whp) if ${\mathbb{P}}(E_{n})\rightarrow 1$ as $n\rightarrow\infty$. we have $\zeta({G_{n,1/2}})\sim\chi({G_{n,1/2}})\sim\frac{n}{2\log_{2}n}$.
Erdős and Gimbel [4] (see also [5]) asked the following question: For $G\sim G_{n,1/2}$, does the difference $\chi(G)-\zeta(G)$ tend to infinity as $n\rightarrow\infty$? In other words, is there a function $f(n)\rightarrow\infty$ such that, with high probability,
$\chi(G)-\zeta(G)>f(n)?$
At a conference on random graphs in Poznań222most likely in 1991, or possibly in 1989, based on Erdős’s and Gimbel’s participation records, Erdős offered \$100 for the solution if the answer was ‘yes’, and \$1000 if the answer was ‘no’ (although later said to Gimbel that perhaps \$1000 was too much) [5]. The question is listed as Problem #625 on Thomas Bloom’s Erdős Problems website [1].
In this note, we show that it is not the case that the difference $\chi(G)-\zeta(G)$ is whp bounded, and that in fact, it is not whp bounded by $n^{1/2-o(1)}$. It turns out that any function $g(n)$ so that whp $\chi(G)-\zeta(G)\leqslant g(n)$ cannot be smaller than the concentration interval length of the chromatic number $\chi({G_{n,1/2}})$, for which corresponding lower bounds were recently obtained [6, 8, 7]. Formally we prove the following statement.
Theorem 1.
Let $G\sim G_{n,1/2}$. There is a constant $c>0$ so that for any sequence of integers $g(n)$ such that
${\mathbf{P}}\big{(}\chi(G)-\zeta(G)\leqslant g(n)\big{)}>0.999,$ (1)
there is a sequence of integers $n^{*}$ such that
$g(n^{*})>c\frac{\sqrt{n^{*}}\log\log n^{*}}{\log^{3}n^{*}}.$
The proof relies on a comparison of the chromatic and co-chromatic numbers of $G\sim G_{n,1/2}$ and of the complement graph $\bar{G}\sim G_{n,1/2}$, which is obtained from $G$ by exchanging all the edges and non-edges. Clearly $\zeta(G)=\zeta(\bar{G})$, so for a function $g(n)$ as above, $\chi(G)$ and $\chi(\bar{G})$ are likely to be at most $g(n)$ apart from each other. Of course $\chi(G)$ and $\chi(\bar{G})$ have the same distribution $\chi({G_{n,1/2}})$. Informally speaking, to see why $g(n)$ has to be at least the concentration interval length of this distribution, suppose that we have some lower bound on the concentration interval length; for example, a statement saying that any interval containing $\chi({G_{n,1/2}})$ with probability at least 0.9 has length at least $\ell(n)$. Then if $X_{1},X_{2}\sim\chi(G_{n,1/2})$ were independent samples of this distribution, they would be reasonably likely to be at least about $\ell(n)$ apart from each other. Of course $\chi(G)$ and $\chi(\bar{G})$ are not independent, but the first is an increasing and the other a decreasing function of the edges of $G\sim{G_{n,1/2}}$, and so with the help of Harris’s Lemma we can draw the same conclusion. So we know that $\chi(G)$ and $\chi(\bar{G})$ are both likely to be at most $g(n)$ apart, but also reasonably likely to be at least $\ell(n)$ apart and it follows that $g(n)\geqslant\ell(n)$.
Independently from this note, Raphael Steiner recently also discovered the connection between $\chi(G)-\zeta(G)$ and the concentration interval length of $\chi(G_{n,1/2})$. For his work on this and two other questions of Erdős, Gimbel and Straight, see [10].
2 Proof of Theorem 1
Turning to the details, let us first state the non-concentration result for $\chi({G_{n,1/2}})$ that we will use, which follows by combining Theorem 8 from [8] and Theorem 1.2 from [7].
Theorem 2 ([8, 7]).
There is a constant $c>0$ so that for any sequence of intervals $[s_{n},t_{n}]$ such that ${\mathbf{P}}\big{(}\chi(G_{n,1/2})\in[s_{n},t_{n}])>0.9$, there is a sequence of integers $n^{*}$ such that
$t_{n^{*}}-s_{n^{*}}>c\frac{\sqrt{n^{*}}\log\log n^{*}}{\log^{3}n^{*}}.$
Theorem 1 follows directly from Theorem 2 and the following proposition.
Proposition 3.
Let $g(n)$ be a sequence of integers which satisfy (1), then there is a sequence of intervals $[s_{n},t_{n}]$ with $t_{n}-s_{n}=g(n)$ so that
${\mathbf{P}}\big{(}\chi(G_{n,1/2})\in[s_{n},t_{n}])>0.9.$
Proof of Proposition 3.
Let $G\sim{G_{n,1/2}}$ and let $\bar{G}$ be the complement graph of $G$ (which contains exactly the edges which are missing in $G$). Then $\bar{G}\sim{G_{n,1/2}}$ and $\zeta(\bar{G})=\zeta(G)$, so with probability at least 0.999,
$\chi(\bar{G})\leqslant\zeta(\bar{G})+g(n)=\zeta(G)+g(n)\leqslant\chi(G)+g(n).$ (2)
Now let $s_{n}$ be the smallest integer $k$ such that ${\mathbf{P}}(\chi(G)\leqslant k)\geqslant 0.05$, and define the following events:
$\mathcal{D}$ =$=\{\chi(G)\leqslant s_{n}\},$
$\mathcal{U}$ =$=\{\chi(\bar{G})\leqslant s_{n}+g(n)\}.$
Then $\mathcal{D}$ is a down-set and $\mathcal{U}$ is an up-set in the edges of $G\sim{G_{n,1/2}}$. By the definition of $s_{n}$,
${\mathbf{P}}(\mathcal{D})\geqslant 0.05.$
Furthermore, if $\mathcal{D}$ holds, then either (2) does not hold (which has probability at most 0.001), or (2) holds and implies $\mathcal{U}$, so
${\mathbf{P}}(\mathcal{U}\cap\mathcal{D})\geqslant{\mathbf{P}}(\mathcal{D})-{\mathbf{P}}\big{(}\chi(\bar{G})>\chi(G)+g(n)\big{)}\geqslant{\mathbf{P}}(\mathcal{D})-0.001.$
By Harris’s Lemma (see for example §2, Lemma 3 in [3]),
${\mathbf{P}}(\mathcal{U})\geqslant{\mathbf{P}}(\mathcal{U}\cap\mathcal{D})/{\mathbf{P}}(\mathcal{D})\geqslant 1-\frac{0.001}{{\mathbf{P}}(\mathcal{D})}\geqslant 1-\frac{0.001}{0.05}=0.98.$
But since $G$ and $\bar{G}$ have the same distribution, and by the definition of $s_{n}$, this implies
${\mathbf{P}}\big{(}s_{n}\leqslant\chi(G)\leqslant s_{n}+g(n)\big{)}$ $=1-{\mathbf{P}}\big{(}\chi(G)\leqslant s_{n}-1\big{)}-{\mathbf{P}}\big{(}\chi(G)>s_{n}+g(n)\big{)}$
$\geqslant 1-0.05-0.02>0.9.$
The claim follows. ∎
3 Discussion
So how about Erdős and Gimbel’s original question: does $\chi(G)-\zeta(G)$ tend to infinity whp for $G\sim{G_{n,1/2}}$? Theorem 1 suggests that the answer is ‘yes’, but of course does not imply this.
If we had a result like Theorem 1, but with the conclusion that $g(n)\geqslant h(n)$ for every $n$ and some $h(n)\gg\sqrt{n}/\log n$, this would imply that the answer to the original question is ‘yes’: by an argument of Alon [2, 9], both $\chi({G_{n,1/2}})$ and $\zeta({G_{n,1/2}})$ are contained in a sequence of intervals of length about $\sqrt{n}/\log n$, respectively, and consequently so is their difference. [^3] Formally the statement is: For $G\sim G_{n,1/2}$ and any function $\omega(n)\rightarrow\infty$, there is a sequence of intervals of length at most $\omega(n)\sqrt{n}/\log n$ that contains $\chi(G)-\zeta(G)$ whp. Let $[s_{n},t_{n}]$ be such a sequence of intervals, then taking $g(n)=t_{n}$ would give that $t_{n}\geqslant h(n)\gg\sqrt{n}/\log n$, which implies that $s_{n}\gg\sqrt{n}/\log n$ as well444for a suitable choice of the arbitrary function $\omega(n)\rightarrow\infty$, and so whp $\chi(G)-\zeta(G)\geqslant s_{n}\gg\sqrt{n}/\log n$.
It is reasonable to expect that the chromatic number $\chi({G_{n,1/2}})$ is close to its first moment threshold, that is, the smallest $k$ such that the expected number of $k$-colourings is at least $1$. [^5] This is an oversimplification and there are several complications with this heuristic, for a detailed discussion see §1.3 and the appendix in [8]. Using the same heuristic for the cochromatic number $\zeta({G_{n,1/2}})$, the first moment threshold there should be of order $n/\log^{3}n$ smaller than that of the chromatic number: for any $k\sim n/(2\log_{2}n)$, the expected number of $k$-cocolourings is multiplied by a factor $2^{k}=\exp(\Theta(n/\log n))$ when compared to the expected number of $k$-colourings (since we may choose for each colour class whether it is a clique or an independent set); and decreasing the number of colours by 1 should multiply the expectation by a factor $\exp(-\Theta(\log^{2}n))$ (like it does for the chromatic number [^6] Again an oversimplification, and we refer to the discussion in [8].). We therefore make the following conjecture.
Conjecture 4.
For $G\sim{G_{n,1/2}}$, whp,
$\chi(G)-\zeta(G)=\Theta(n/\log^{3}n).$
Acknowledgements
I would like to thank John Gimbel for bringing this problem to my attention and for helpful discussions, and Svante Janson for his help in tracking down past conference proceedings.